OpenStudy (anonymous):

Is one more correct than the other? \[\int\limits_{a}^{t} f'(x) = f(t) - f(a)\] \[\int_{a}^{t} f(x) = F(t) - F(a) \]

OpenStudy (anonymous):

\[\int\limits_{a}^{t} f'(x) = f(t) - f(a)\] \[\int_{a}^{t} f(x) = F(t) - F(a) \]

OpenStudy (irishboy123):

"." if it is wrt dx, they are the same

OpenStudy (anonymous):

oh yes.. missing the dx So Either would be acceptable if I was using as an explanation for my work?

OpenStudy (anonymous):

\[\int\limits_{a}^{t} f'(x) dx = f(t) - f(a)\] \[\int\limits_{a}^{t} f(x) dx = F(t) - F(a)\]

OpenStudy (irishboy123):

@oldrin.bataku

OpenStudy (irishboy123):

yep

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

I prefer the second